# Improving Math Problem Solving in Large Language Models Through Categorization and Strategy Tailoring

### v2

@@ Abstract
In this paper, we investigate how to harness large language models (LLMs) to solve mathematical problems both quickly and accurately. Specifically, we demonstrate the effectiveness of classifying problems into distinct categories and applying category-specific problem-solving strategies to enhance the math performance of LLMs. We develop a straightforward machine learning model for problem categorization and show that its accuracy can be significantly improved through the creation of well-designed training datasets. We believe that our approach works by helping reduce hallucinations in LLMs, which is a critical step toward unlocking their potential to tackle advanced mathematical problems.
@@ 1 Introduction
Unsurprisingly, many AI models struggle with mathematical problems, frequently producing inaccurate results. For instance, even widely known models like ChatGPT often fail to solve simple competition math problems correctly [8]. One primary reason is hallucination—where the model generates responses based on faulty logic or incorrect assumptions.
@@
To address these challenges, researchers at Google DeepMind have developed models like AlphaGeometry and AlphaProof, which integrate text generation with formal reasoning and theorem-proving techniques [13]. These models have shown promising results: for example, AlphaProof achieved a score of 28 on the 2024 International Mathematics Olympiad (IMO), earning a silver medal [18], while AlphaGeometry outperformed the average IMO silver medalist on geometry problems [16]. However, these models have limitations. AlphaProof, for instance, can take up to three days to solve some IMO problems [18], far exceeding competition time limits and making it impractical for real-time solutions. Additionally, AlphaGeometry is restricted to geometry problems, and AlphaProof performs poorly on combinatorics tasks [19].
@@
This paper focuses on the specific but significant challenge of enabling LLMs to compute solutions to mathematical problems both quickly and accurately. Unlike AlphaProof and AlphaGeometry, our approach targets computational problems that require producing numeric answers. Even within this narrower scope, hallucinations in LLM outputs can lead to particularly egregious errors.
@@
We propose a novel approach that helps mitigate hallucination by categorizing problems into predefined categories and providing category-specific solving strategies as input to the LLM. Our framework identifies four primary categories—algebra, combinatorics, geometry, and number theory—but can be extended to accommodate more nuanced categorizations. Depending on the problem’s category, we employ one of two problem-solving strategies: chain of thought or program of thought. This structured methodology substantially improves the accuracy of LLM problem-solving.
@@
To automate problem categorization, we developed a lightweight deep neural network model. The key to its effectiveness lies in curating the right training data. In particular, we show the importance of domain-specific factors such as “answer extraction” and achieving a good balance of different problem categories. We then show how to probabilistically associate a problem-solving strategy with each identified category.
@@
Our empirical analysis demonstrates that our categorization model achieves over 80% accuracy. Furthermore, we evaluate the impact of our neural model-categorization-driven strategy selection against two baselines: using ground-truth categories for strategy selection and randomly assigning strategies. Our approach outperforms random strategy assignment by 67% but falls 29% short of the performance achieved using ground-truth categories.
@@
    We develop a lightweight neural model for problem categorization.
@@
    We show the importance of - and the method for - obtaining the right training data for effective neural categorization.
@@ 2 Background
Hallucination occurs when an LLM produces reasoning that is incorrect or generates information that is not necessarily true. This issue is particularly critical in mathematical problem-solving. If an LLM hallucinates and arrives at an incorrect intermediate result inconsistent with the given information, it is highly unlikely to recover and successfully solve the problem.
Our overall approach enhances LLMs’ reliability by introducing category-specific math problem-solving strategies. These strategies help mitigate hallucinations by providing the model with more structured and precise prompts, offering greater context to guide the solution process. In particular, we use two strategies: Chain of Thought (CT) [17] and Program of Thought (PT) [5].
Chain of Thought [17] involves asking an LLM to come up with a detailed, step-by-step solution to a problem before returning the final answer. This approach encourages logical reasoning, which helps reduce errors and hallucinations in problem-solving.
@@
While CT can be applied broadly to any mathematical problem to mitigate hallucination, it is particularly effective for problems requiring careful reasoning. On the other hand, PT is better suited for problems involving iteration or recursion, as it can execute such operations efficiently and accurately.
@@ 5 Categorization Model
One very effective way to predict the category of a problem is to use word frequency. This is because some words can be considered ‘indicators’, or words that indicate that a problem is very likely to be of a particular . Table 2 shows examples of several indicators found in the MATH dataset:
@@
This model achieves between a $60\%$ and $70\%$ accuracy with the categorization of problems around the difficulty level of the AIME (see Section 7 for suitable results). We now examine how to further improve the model’s accuracy.
@@ 7.1.1 On Further Improving Categorization
Here are the problems that both models got wrong during the categorization test:
@@ 7.2 Downstream Task Evaluation
    The model answers 20% of the problems correctly, and has a nonzero correct answer frequency on 32% of the problems. These values are slightly worse than the values for perfect categorization (28%, 32%), and significantly better than the values for random choice (12%, 24%).
@@ 8 Conclusion
We find that the accuracy of LLMs when a strategy is chosen using perfect categorization is much higher than the accuracy of LLMs when a random strategy is chosen. The accuracy of LLMs when a strategy is chosen using our categorization model lies in between these two.
Our model achieves around 84% accuracy in predicting the categories of problems. While this is fairly high, the accuracy can likely be further improved by using a larger and even more accurate dataset.

### v3

@@ Abstract
In this paper, we explore how to leverage large language models (LLMs) to solve mathematical problems efficiently and accurately. Specifically, we demonstrate the effectiveness of classifying problems into distinct categories and employing category-specific problem-solving strategies to improve the mathematical performance of LLMs. We design a simple yet intuitive machine learning model for problem categorization and show that its accuracy can be significantly enhanced through the development of well-curated training datasets. Additionally, we find that the performance of this simple model approaches that of state-of-the-art (SOTA) models for categorization. Moreover, the accuracy of SOTA models also benefits from the use of improved training data. Finally, we assess the advantages of using category-specific strategies when prompting LLMs and observe significantly better performance compared to non-tailored approaches.
@@ 1 Introduction
Unsurprisingly, many AI models struggle with mathematical problems, frequently producing inaccurate results. For instance, even widely known models like ChatGPT often fail to solve simple competition math problems correctly [9, 13]. It has been observed that many AI models generate responses based on faulty logic or incorrect assumptions.
@@
To address these challenges, researchers at Google DeepMind have developed models like AlphaGeometry [18] and AlphaProof [17], which integrate text generation with formal reasoning and theorem-proving techniques. These models have shown promising results: for example, AlphaProof achieved a score of 28 on the 2024 International Mathematics Olympiad (IMO), earning a silver medal, while AlphaGeometry outperformed the average IMO silver medalist on geometry problems. However, these models have limitations. AlphaProof, for instance, can take up to three days to solve some IMO problems, far exceeding competition time limits and making it impractical for real-time solutions. Additionally, AlphaGeometry is restricted to geometry problems, and AlphaProof performs poorly on combinatorics tasks.
@@
This paper focuses on the specific but significant challenge of enabling LLMs to compute solutions to mathematical problems both quickly and accurately. Unlike AlphaProof and AlphaGeometry, our approach targets computational problems that require producing numeric answers. Even within this narrower scope, LLM outputs can have particularly egregious errors.
@@
We propose a novel approach that helps improve LLMs’ math-solving ability by categorizing problems into predefined categories and providing category-specific solving strategies as input to the LLM. Our framework identifies four primary categories—algebra, combinatorics, geometry, and number theory—but can be extended to accommodate more nuanced categorizations. Depending on the problem’s category, we employ one of two problem-solving strategies: chain of thought or program of thought. This structured methodology substantially improves the accuracy of LLM problem-solving.
@@
To automate problem categorization, we developed a lightweight deep neural network model. The key to its effectiveness lies in curating the right training data. In particular, we show the importance of domain-specific factors such as “answer extraction” and achieving a good balance of different problem categories. We then show how to probabilistically associate a problem-solving strategy with each identified category. We also compare our simple neural model with a more sophisticated approach that leverages SOTA models (ChatGPT’s GPT-4o) with in-context learning [7] for problem categorization.
@@
Our empirical analysis demonstrates that our simple categorization model achieves 84% accuracy, which is close to that of SOTA models (92%). Furthermore, we evaluate the impact of our neural model-categorization-driven strategy selection against two baselines: using ground-truth categories for strategy selection and randomly assigning strategies. Our approach outperforms random strategy assignment by 67% but falls 29% short of the performance achieved using ground-truth categories.
@@
    We develop a lightweight neural model for problem categorization and compare it with sophisticated and heavy-weight SOTA models.
@@
    We show the importance of - and the method for - obtaining the right training data for effective neural categorization. We find that this insight holds for both our simple model and SOTA models.
@@ 2 Background
LLMs can produce reasoning that is often incorrect or generates information that is not necessarily true. This issue is particularly critical in mathematical problem-solving. If an LLM arrives at an incorrect intermediate result inconsistent with the given information, it is highly unlikely to recover and successfully solve the problem.
Our overall approach enhances LLMs’ reliability by introducing category-specific math problem-solving strategies. These strategies help LLMs by providing them with more structured and precise prompts, offering greater context to guide the solution process. In particular, we use two strategies: Chain of Thought (CT) [19] and Program of Thought (PT) [5].
Chain of Thought [19] involves asking an LLM to come up with a detailed, step-by-step solution to a problem before returning the final answer. This approach encourages logical reasoning, which helps reduce errors and incorrect reasoning in problem-solving.
@@
While CT can be applied broadly to any mathematical problem to mitigate erroneous reasoning, it is particularly effective for problems requiring careful reasoning. On the other hand, PT is better suited for problems involving iteration or recursion, as it can execute such operations efficiently and accurately.
@@ 5 A Simple Categorization Model
One very effective way to predict the category of a problem is to use word frequency. This is because some words can be considered ‘indicators’, or words that indicate that a problem is very likely to be of a particular. Table 2 shows examples of several indicators found in the MATH dataset. In our approach, we pick all words that have length at least 3 characters and appear at least 5 times in the training dataset as indicators.
@@
To extract features from a problem, we look at whether the word corresponding to each indicator is in the problem or not. If the word is in the problem, the feature is set to $1$, and otherwise it is set to 00.
@@
This model achieves between a $60\%$ and $70\%$ accuracy with the categorization of problems around the difficulty level of the AIME (see Section 7 for suitable results).
5.1 A More Sophisticated Categorization Model?
Given the simplicity of the model described above, one might wonder how a more complex model would perform in problem categorization. To explore this, we utilize ChatGPT-4o, a highly advanced large language model (LLM). Specifically, we employ in-context learning [7], where we prompt ChatGPT with a series of problem-category pairs from our training set (serving as “demonstrations”), followed by asking it to categorize a series of problems from our test set.
Our prompt is structured as below:
I am going to give you the task of categorizing mathematics competition problems into four categories. The four categories are “algebra”, “combinatorics”, “geometry”, and “number theory”. I will give you several example problems drawn from recent mathematics competitions, along with their categories. Based on this, I would like you to categorize new problems.
$<$a series of ”Problem:”-”Category:” pairs drawn from our training data$>$
The above 25 problems have been categorized into four categories: “algebra”, “geometry”, “combinatorics”, and “number theory”. Based on the above, please categorize each of the following 10 problems into either “algebra”, “geometry”, “combinatorics”, or “number theory”. Your answer should be formatted like the above.
$<$a series of ”Problem”s drawn from our test set$>$
As expected, this model out-performs our simple model, achieving around 88% accuracy in problem categorization (further discussion in Section 7).
We now examine how to further improve our problem categorization accuracy.
@@ 7.1.1 On Further Improving Categorization
In this section, we will explore discrepancies that still exist in our updated model. Here are the problems that both our initial and improved models got wrong during the categorization test:
@@
7.1.2 Performance of a More Sophisticated Categorization Model
We now conduct a small-scale evaluation to compare our simple categorization model that uses a three-layer neural network with a more sophisticated model that uses ChatGPT in-context learning. We conduct two tests to study performance before and after adding AIME problems to ChatGPT’s in-context learning.
In the first test, we include 25 problem-category pairs in our prompt to ChatGPT, all drawn from the MATH dataset. These pairs consist of 7 algebra problems, 6 combinatorics problems, 6 geometry problems, and 6 number theory problems. In the second test, we again use 25 problem-category pairs in the prompt. However, only 13 of these problems are from the MATH dataset, while the remaining 12 are sourced from the AIME dataset.
In both tests, we then ask ChatGPT to categorize 25 AIME-like problems (the same ones which are used in Tables 3 and 4.
We make two main observations:
  1. 1.
    ChatGPT gets $22$ categorizations out of $25$ correct when using the MATH dataset, whereas it gets $23$ categorizations correct when using a training dataset composed of both MATH dataset problems and AIME problems. This shows that using the correct data can even help large SOTA models’ problem categorization.
  2. 2.
    ChatGPT’s accuracy when using the MATH dataset is $88\%$, which is only slightly better than than our updated model’s total accuracy ($84\%$). This shows that when given the right data, even small models like the one we developed are sufficiently competitive for problem categorization.
@@ 7.2 Downstream Task Evaluation
    The model answers 20% of the problems correctly, and has a nonzero correct answer frequency on 32% of the problems. These values are slightly worse than the values for perfect categorization (28%, 36%), and significantly better than the values for random choice (12%, 24%).
@@ 8 Conclusion
We find that the accuracy of LLMs when a strategy is chosen using perfect categorization is much higher than the accuracy of LLMs when a random strategy is chosen. The accuracy of LLMs, when a strategy is chosen using our simple categorization model, lies in between these two.
Our simple neural model achieves around 84% accuracy in predicting the categories of problems, which we find to be comparable to that achievable by SOTA models (such as ChatGPT with in-context learning). This high accuracy is helped in particular by well-curated training data, and can be improved even further by using larger datasets with similar rich examples.